two consecutive integers are such that three times the first integer is 15 more than the second integer. find the integers
Answer by ReadingBoosters(3246)

x x+1
...
3x = 15+x+1
3x = 16 +x
3x - x = 16
2x = 16
x = 8
...
Integers are 8 and 9
...
Check
3(8) = 24
15 + 8 + 1 = 24
.....................
Delighted to help; feel free to email if you have further questions.
HomeworkHelpers@readingboosters.com
-Reading Boosters
Website: www.MyHomeworkAnswers.com
Wanting for others what we want for ourselves.